# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v1

@@ I.1 The Tsirelson inequality
In 2006, Boris Tsirelson wrote a paper with a title that asks an endearingly straightforward question: “How often is the coordinate of a harmonic oscillator positive?” [1]. Tsirelson demonstrated that a simple dynamical assumption about classical oscillators can give rise to a surprising disparity between classical and quantum expectations. Violations of the “Tsirelson inequality” [2] are associated inherently with the negativity of the Wigner function [3], and present an alluring scheme to assess the “quantumness” of a system using only single-time measurements on an ensemble of systems.
This work went virtually unnoticed for more than a decade and a half, until it was rediscovered by Zaw et al., who in their 2022 paper extended Tsirelson’s work to a broader family of systems undergoing uniform precession, such as finite-dimensional spins in real space [4]. Tsirelson’s protocol has been generalised as a viable entanglement witness for coupled harmonic oscillator systems [5, 6] and entangled spin ensembles [7], and the most recent work in the field has focussed primarily upon further generalisations of the protocol [8] and computing stricter bounds on the maximal quantum violations of Tsirelson’s inequality [9, 10]. A recent triumph has been the first experimental realisation (a so-called quantumness certification) of Tsirelson violations in precessing nuclear spins [11].
@@
Thus, the Tsirelson quantity $\langle A\rangle$ can be measured using only single-time measurements on an ensemble of identical systems. In practice, measuring the dichotomic variable $Q$, without ascertaining any further information about the particle’s position, may be done using a light beam that illuminates one half of the harmonic well [12].
However, a crucial point which has remained hitherto unaddressed in the literature is the validity of the dynamical assumption of uniform precession in a quantum-mechanical context. If truly the Tsirelson test is to be interpreted as a credible certification of quantum behaviour, it is imperative to demonstrate that violations are indicative of genuine quantum phenomena and not of simply the failure of its founding dynamical assumption.
Figure 2: Classical phase-space depiction of the system’s dynamics for three possible measurement times. If the system is initialised in the green region for one of the experimental runs, then it will have a probability of $2/3$ of being found on the right side of the origin at a randomly-chosen measurement time $t\in{0,T/3,2T/3}$. Conversely, the probability will be $1/3$ if the particle begins in a blue region. As such, the ensemble average, a convex sum of these probabilities, is bounded between $\frac{1}{3}$ and $\frac{2}{3}$. This is equivalent to Eq. 1.1 up to a scale factor. (Adapted from Zaw et al. [13].)
Uniform precession can be defined perfectly well in the case of a classical oscillator by considering, for instance, the phase-space representation of its trajectory. The condition for uniform precession is then simply that the angular variable $\varphi=\tan^{-1}(x/p)$ evolve linearly in time ($\dot{\varphi}=\text{const.}$). However, quantum-mechanically, the corresponding operator, related to the arrival time operator, is conjugate to the Hamiltonian and is not self-adjoint [14, 15, 16, 17]. As such, the eigenstates of the angular operator cannot be orthogonal, and there is some intrinsic indefiniteness in the angular behaviour and hence in the definition of uniform precession. This property is also related to the difficulties of defining a phase operator for the harmonic oscillator [18]. In practice, this means that uniform precession can be defined in multiple inequivalent ways in quantum mechanics, as we shall see.
@@
Although the Wigner function evolves along the classical path, it is not a probability distribution, since it can go negative, and thus we cannot assert that the evolution in the quantum case respects uniform precession. In the Heisenberg picture with a time independent Wigner function, the Tsirelson quantity is
@@
which makes it clear that Tsirelson inequality violations are a negative Wigner function effect.
We dedicate the remainder of this paper to finding precise sets of conditions for uniform precession, in terms of measurable probabilities, that are all satisfied classically, but which lead to quantum-mechanically inequivalent expressions and measures, which we may use to assess the degree to which the dynamical assumption is satisfied. In order to do so, we will embed the Tsirelson scheme within the more comprehensive Leggett–Garg framework [24]—the canonical test of non-classicality in temporal correlations—and utilise some of its powerful machinery.
@@ I.2 Leggett–Garg tests of non-classicality
The Leggett–Garg inequalities (LGIs) [24, 25], a temporal analogue to Bell-type inequalities [26, 27], must be satisfied by any theory obeying “macrorealism”, a quintessentially classical world-view that is encapsulated by the following three postulates: macroscopic realism per se—a macroscopic system exists in a definite state at all times; non-invasive measurability—the state of the system can be determined, in principle, without affecting its subsequent dynamics; and induction—the state does not depend upon future measurements.
Macrorealism pertains to measurements made on a specific data set, which we take here to be the set of three single-time averages $\langle Q_{i}\rangle$, where $Q_{i}\equiv Q(t_{i})$, and the three correlators $C_{ij}$, the latter measured in a non-invasive way. With possible measurement outcomes $s_{i}=\pm 1$, where $i=1,2,3$ for measurements taken at times $t_{1},t_{2},t_{3}$, respectively, the two-time correlators are given by
@@
where $D$ is the triple correlator, $\langle Q_{1}Q_{2}Q_{3}\rangle$, but is never measured in the simplest protocols. These two quasi-probabilities are classically positive. Quantum-mechanically, they are equivalent to the expressions
@@
Turning now to the conditions for macrorealism, the three postulates of macrorealism imply that there exists an underlying joint probability for the given data set, which implies that a set of twelve two-time and four three-time LG inequalities hold. Specifically, the twelve quasi-probabilities Eq. 1.6—for time pairs $(t_{1},t_{2})$, $(t_{1},t_{2})$, $(t_{1},t_{2})$—are non-negative (LG2s) and the following four LG3 inequalities are satisfied:
@@
These sixteen conditions are not only necessary for macrorealism but also sufficient, as the LG version of Fine’s theorem shows [30, 31, 32]. In quantum mechanics, the LG2 and LG3 inequalities can be violated up to the value $\frac{1}{8}$, the Lüders bound. [^1] This is for so-called Lüders measurements, which simply measure the value of $Q$. More negative values can be obtained by measuring finer-grained quantities and then coarse-graining the probabilities to find probabilities of $Q$ [33, 34].
We also note that the four LG3s may be written as
@@
This means that if the LG3s are satisfied but the LG2s are not, then even though MR is violated for the full data set, it can still be satisfied for the data subset consisting of the sign-change variables, $\chi_{1}=s_{1}s_{2}$ and $\chi_{2}=s_{2}s_{3}$, which are $\pm 1$ depending on whether or not the sign of $Q$ has flipped between measurement times. This is a point we will make extensive use of in Section III, where we are interested in the sign-change probabilities as a quantification of uniform precession.
Violations of the LGIs have been observed in a myriad of physical systems, such as superconducting qubits [35, 36], nuclear spins [37, 38, 39], millimetre-scale crystals [40], photons [41], and (recently) neutron interferometry [42, 43]. The Leggett–Garg inequalities also have far-reaching applications to other fields [44, 45, 46, 47, 48]. A more exhaustive account of the experimental tests of macrorealism to date is provided in the extensive review articles Refs. [33, 49].
@@
But does a violation truly imply genuine quantum behaviour and the failure of macrorealism, or could a stubborn macrorealist now argue that it stems purely from faulty dynamics? A classical system that breaks the dynamical assumption—that is, one which is not precessing uniformly (or not oscillating at all)—will easily violate the Tsirelson inequality. Fig. 3 depicts qualitatively that a Tsirelson violation could be caused by a classical particle that does not obey the expected Newtonian dynamics of oscillation induced by a harmonic potential, lingering too long on one side of the origin; or perhaps we have even simply defined the origin inaccurately. Such a trajectory still has a defined position at all times, and hence is macrorealistic, yet a spurious Tsirelson inequality violation will ensue. So, just as separating realism from non–invasive measurability has proved to be the greatest hassle in LG tests, do we now run into a similar problem in Tsirelson tests of separating the failure of realism from this newfound dynamical assumption?
Figure 3: Unlike the blue classical trajectory, which follows the expected sinusoidal dynamics, the erractic red trajectory, still macrorealistic, dwells for too long on the right side of the origin and will violate the Tsirelson inequality, despite there being no quantum behaviour. This is the faulty dynamics loophole.
Our strategy to assess the dynamical assumption will be to re-write the Tsirelson quantity in terms of various probabilities for two and three time measurements that can be interpreted as measures of uniform precession. We will do this in two ways. Our first method involves demonstrating an explicit connection between $\langle A\rangle$ and the probability of zero sign changes, $L_{1}$ (Eqs. 1.12 and 1.16), which should be zero classically. Our second method stems from the interesting observation that a weaker condition of setting merely the difference q(+,+,+)q(,,)q(+,+,+)-q(-,-,-) to be zero also leads to the Tsirelson inequality, as we shall show, and thus may also be considered a measure of uniform precession.
@@ I.3 This paper
Assessing how well the dynamical assumption of uniform precession holds and classifying the nature of Tsirelson violations hereinafter forms the core narrative of this paper. In Section II, we lay forth the mathematical machinery required for quantum-mechanical calculations and discuss quantum states and their violations of the Tsirelson inequality, proving some results relating to maximally-violating states through a group theoretic perspective. We present our most important results in Section III on the main protocol we have devised to assess uniform precession, broken down into the two methods of defining uniform precession discussed prior. Here we shall require the breadth of the Leggett–Garg formalism to disentangle violations of macrorealism from the breaking of uniform precession. In Section IV, we investigate two other quantities that are less strictly related to the Tsirelson inequality but provide further heuristic evidence of the extent to which uniform precession is satisfied. We will show that one has quantum properties identical to the classical case and the other approximately so. Moreover, we will argue that these results can be used to rule out alternative classical explanations. Finally, we conclude in Section V with a critical assessment of whether the protocol introduced in Section III ultimately sidesteps the problem of non-invasive measurement. We report lengthy technical results and rigorous mathematical proofs in the appendices.
@@ II.1 Expansion in energy eigenbasis
We can compute the Tsirelson quantity for an arbitrary state of the QHO Hilbert space formed as a superposition of energy eigenstates, from the ground state upwards, up to a finite cut-off:
@@
$\langle\Psi|2\hat{\theta}-1|\Psi\rangle=\sum_{n,k=0}^{N}\langle\Psi|n\rangle\langle n|2\hat{\theta}-1|k\rangle\langle k|\Psi\rangle\\ =\sum_{n,k=0}^{N}c_{n}^{*}Q_{nk}c_{k}\ ,$ (2.3)
@@
$|\Psi,t\rangle=\sum_{n=0}^{N}c_{n}e^{-in\omega t}|n\rangle\ ,$ (2.5)
since $|n\rangle$ is by definition an eigenstate of the QHO Hamiltonian with eigenvalue $\omega(n+\frac{1}{2})$. We can drop the vacuum contribution of $\frac{1}{2}\omega$ since this will have no effect upon the expectation value:
@@
Violations of the Tsirelson inequality are manifestly associated with negativity in the Wigner function. A corollary is that a single coherent state, which has a completely positive Wigner function, will not violate the Tsirelson inequality. Neither will a single energy eigenstate, for it is trivially of strictly odd or even parity; a neat proof, in terms of probability currents, that definite-parity states do not violate the Tsirelson inequality is provided in Section IV.1. Perhaps more surprising is that a superposition of at least three energy eigenstates is required to produce a violation. We provide a full proof of this fact in Appendix E.
@@ II.2 Maximally violating eigenstates
with a upper-bound Tsirelson violation $\langle A\rangle=1.1195$; or equivalently, by swapping the sign of the odd-parity eigenstate,
@@ II.3 Structure of the Tsirelson operator
Because the two normal operators commute, they cannot mix each other’s eigenspaces. Notice that the action of $U$ on the basis states $|n\rangle$ clearly partitions the Hilbert Space, $\mathcal{H}$, into 3 degenerate eigenspaces corresponding to its three eigenvalues $\lambda_{j}$:
@@
Figure 5: The Tsirelson operator in the energy eigenbasis with any finite energy cut-off can be interpreted as the adjacency matrix of a simple weighted graph. This is illustrated with a cut-off at $n=29$, where the nodes are energy eigenstates and the edges represent non-zero matrix elements of the Tsirelson operator (the colour representing the value of the matrix element).
@@ III.1 General Framework
We now come to the main issue, which is assessing the degree to which the dynamical assumptions behind the Tsirelson inequality hold. As discussed, like many other classical phenomena, uniform precession and its closely-related assumptions have no unique definition in quantum theory. We will therefore explore a number of different, inequivalent implementations which entail different measurement techniques to check it. Specifically, we will implement the following procedure in a number of different ways:

### v2

@@ I.1 The Tsirelson inequality
In 2006, Boris Tsirelson wrote a paper with a title that asks an endearingly straightforward question: “How often is the coordinate of a harmonic oscillator positive?” [1]. Tsirelson demonstrated that a simple dynamical assumption about classical oscillators can give rise to a surprising disparity between classical and quantum expectations. Violations of the “Tsirelson inequality” [2] are associated inherently with the negativity of the Wigner function [3] and present an alluring scheme to assess the “quantumness” of a system using only single-time measurements on an ensemble of systems.
This work went virtually unnoticed for more than a decade and a half until it was rediscovered by Zaw et al., who in their 2022 paper extended Tsirelson’s work to a broader family of systems undergoing uniform precession, such as finite-dimensional spins in real space [4]. Tsirelson’s protocol has been generalised as a viable entanglement witness for coupled harmonic oscillator systems [5, 6] and entangled spin ensembles [7], and the most recent work in the field has focussed primarily upon further generalisations of the protocol [8] and computing stricter bounds on the maximal quantum violations of Tsirelson’s inequality [9, 10]. A recent triumph has been the first experimental realisation (a so-called quantumness certification) of Tsirelson violations in precessing nuclear spins [11].
@@
Thus, the Tsirelson quantity $\langle A\rangle$ can be measured using only single-time measurements on an ensemble of identical systems. In practice, measuring the dichotomic variable $Q$ without ascertaining any further information about the particle’s position may be done using a light beam that illuminates one half of the harmonic well [12].
However, a crucial point which has remained hitherto unaddressed in the literature is the validity of the dynamical assumption of uniform precession in a quantum-mechanical context. If truly the Tsirelson test is to be interpreted as a credible certification of quantum behaviour, it is imperative to demonstrate that violations are indicative of genuine quantum phenomena and not simply of the failure of its founding dynamical assumption.
Figure 2: Classical phase-space depiction of the system’s dynamics for three possible measurement times. If the system is initialised in the green region for one of the experimental runs, then it will have a probability of $2/3$ of being found on the right side of the origin at a randomly-chosen measurement time $t\in\{0,T/3,2T/3\}$. Conversely, the probability will be $1/3$ if the particle begins in a blue region. As such, the ensemble average, a convex sum of these probabilities, is bounded between $1/3$ and $2/3$. This is equivalent to Eq. 1.1 up to a scale factor. (Adapted from Zaw et al. [13].)
Uniform precession can be defined perfectly well in the case of a classical oscillator by considering, for instance, the phase-space representation of its trajectory. The condition for uniform precession is then simply that the angular variable $\varphi=\tan^{-1}(x/p)$ evolve linearly in time ($\dot{\varphi}=\text{const.}$). However, quantum-mechanically, the corresponding operator, related to the arrival time operator, is conjugate to the Hamiltonian and is not self-adjoint [14, 15, 16, 17]. As such, the eigenstates of the angular operator cannot be orthogonal, and there is some intrinsic indefiniteness in the angular behaviour and hence in the definition of uniform precession; this property is also related to the difficulties of defining a phase operator for the harmonic oscillator [18]. In practice, this means that uniform precession can be defined in multiple inequivalent ways in quantum mechanics, as we shall see.
@@
Although the Wigner function evolves along the classical path, it is not a probability distribution, since it can go negative, and thus we cannot immediately assert that the evolution in the quantum case respects uniform precession. In the Heisenberg picture, with a time-independent Wigner function, the Tsirelson quantity is
@@
which makes it clear that Tsirelson inequality violations are indeed effected by negativity in the Wigner function.
We dedicate the remainder of this paper to finding precise sets of conditions for uniform precession in terms of measurable probabilities that are all satisfied classically but which lead to quantum-mechanically inequivalent expressions and measures, which we may use to assess the degree to which the dynamical assumption is satisfied. In order to do so, we will embed the Tsirelson scheme within the more comprehensive Leggett–Garg framework [24]—the canonical test of non-classicality in temporal correlations—and utilise some of its powerful machinery.
@@ I.2 Leggett–Garg tests of non-classicality
The Leggett–Garg inequalities (LGIs) [24, 25], a temporal analogue to Bell-type inequalities [26, 27], must be satisfied by any theory obeying “macrorealism”, a quintessentially classical world-view that is encapsulated by the following three postulates: macroscopic realism per se—a macroscopic system exists in a definite state at all times; non-invasive measurability—the state of the system can be determined, in principle, without affecting its subsequent dynamics; and induction—the current state does not depend upon future measurements.
Macrorealism pertains to measurements made on a specific data set [28], which we take here to be the set of three single-time averages $\langle Q_{i}\rangle$, where $Q_{i}\equiv Q(t_{i})$, and the three correlators $C_{ij}$, the latter measured in a non-invasive way. With possible measurement outcomes $s_{i}=\pm 1$, where $i=1,2,3$ for measurements taken at times $t_{1},t_{2},t_{3}$, respectively, the two-time correlators are defined as
@@
where $D=\langle Q_{1}Q_{2}Q_{3}\rangle$ is the triple correlator but is never measured in the simplest protocols. These two quasi-probabilities are classically positive. Quantum-mechanically, they are equivalent to the expressions
@@
Turning now to the conditions for macrorealism, the three postulates of macrorealism imply that there exists an underlying joint probability for the given data set, which implies that a set of twelve two-time and four three-time LG inequalities hold. Specifically, the twelve quasi-probabilities Eq. 1.6, for time pairs $(t_{1},t_{2})$, $(t_{1},t_{2})$, $(t_{1},t_{2})$, are non-negative (LG2s) and the following four LG3 inequalities are satisfied:
@@
These sixteen conditions are not only necessary for macrorealism but are also sufficient, as is shown by the LG version of Fine’s theorem [31, 32, 33]. In quantum mechanics, the LG2 and LG3 inequalities can be violated up to the value $\frac{1}{8}$, the Lüders bound. [^1] This is for so-called Lüders measurements, which simply measure the value of $Q$. More negative values can be obtained by measuring finer-grained quantities and then coarse-graining the probabilities to find probabilities of $Q$ [34, 35].
We also note that the four LG3 quantities may be written as
@@
This means that if the LG3s are satisfied but the LG2s are not, then even though MR is violated for the full data set, it can still be satisfied for the data subset consisting of the sign-change variables $\chi_{1}=s_{1}s_{2}$ and $\chi_{2}=s_{2}s_{3}$, which are $\pm 1$ depending on whether or not the sign of $Q$ has flipped between measurement times. This is a point we will make extensive use of in Section III, where we are interested in the sign-change probabilities as a quantification of uniform precession.
Violations of the LGIs have been observed experimentally in a myriad of physical systems, such as superconducting qubits [36, 37], nuclear spins [38, 39, 40], millimetre-scale crystals [41], photons [42], and (recently) neutron interferometry [43, 44]. The Leggett–Garg inequalities also have far-reaching applications to other fields [45, 46, 47, 48, 49]. A more exhaustive account of the experimental tests of macrorealism to date is provided in the extensive review articles Refs. [34, 50].
@@
But does a violation truly imply genuine quantum behaviour and the failure of macrorealism, or could a stubborn macrorealist now argue that it stems purely from faulty dynamics? A classical system that breaks the dynamical assumption—that is, one which is not precessing uniformly (or not oscillating at all, even)—will easily violate the Tsirelson inequality. Fig. 3 depicts qualitatively that a Tsirelson violation could be caused by a classical particle that does not obey the expected Newtonian dynamics of oscillation induced by a harmonic potential, lingering for too long on one side of the origin; or perhaps we have just defined the origin inaccurately. Either way, such a trajectory still has a defined position at all times, and hence is macrorealistic, yet a spurious Tsirelson inequality violation will ensue. So, just as separating realism from non–invasive measurability has proved to be the greatest hassle in LG tests, do we now run into a similar problem in Tsirelson tests of separating the failure of realism from this newfound dynamical assumption?
Figure 3: Unlike the blue classical trajectory, which follows the expected sinusoidal dynamics, the erratic red trajectory, still macrorealistic, dwells for too long on the right side of the origin and will violate the Tsirelson inequality, despite there being no quantum behaviour. This is the faulty dynamics loophole.
Our strategy to assess the dynamical assumption will be to re-write the Tsirelson quantity in terms of various probabilities for two-time and three-time measurements that can be interpreted as measures of uniform precession. We will do this in two ways. Our first method involves demonstrating an explicit connection between $\langle A\rangle$ and the probability of zero sign changes, $L_{1}$ (Eqs. 1.12 and 1.16), which should be zero classically. Our second method stems from the interesting observation that a weaker condition of setting merely the difference q(+,+,+)q(,,)q(+,+,+)-q(-,-,-) to be zero also leads to the Tsirelson inequality, as we shall show, and thus may also be considered a measure of uniform precession.
@@ I.3 This paper
Assessing how well the dynamical assumption of uniform precession holds and classifying the nature of Tsirelson violations hereinafter forms the core narrative of this paper. In Section II, we lay forth the mathematical machinery required for quantum-mechanical calculations and discuss quantum states and their violations of the Tsirelson inequality, proving some results relating to maximally-violating states through a group theoretic perspective. We present our most important results in Section III on the main protocol we have devised to assess uniform precession, broken down into the two methods discussed prior. Here we shall require the breadth of the Leggett–Garg formalism to disentangle violations of macrorealism from the breaking of uniform precession. In Section IV, we investigate two other quantities that are less strictly related to the Tsirelson inequality but provide further heuristic evidence of the extent to which uniform precession is satisfied. We will show that one has quantum properties identical to the classical case and the other approximately so. Moreover, we will argue that these results can be used to rule out alternative classical explanations. Finally, we conclude in Section V with a critical assessment of whether the protocol introduced in Section III ultimately sidesteps the problem of non-invasive measurement. We report lengthy technical results and mathematical proofs in the appendices.
@@ II.1 Expansion in energy eigenbasis
We can compute the Tsirelson quantity for an arbitrary state of the QHO Hilbert space formed as a superposition of energy eigenstates from the ground state upwards up to a finite cut-off:
@@
$\langle\Psi|2\hat{\theta}-1|\Psi\rangle=\sum_{n,k=0}^{N}c_{n}^{*}Q_{nk}c_{k}\ ,$ (2.3)
@@
$|\Psi,t\rangle=\sum_{n=0}^{N}c_{n}e^{-int}|n\rangle\ ,$ (2.5)
since $|n\rangle$ is by definition an eigenstate of the QHO Hamiltonian with eigenvalue $(n+\frac{1}{2})$. We can drop the vacuum contribution of $\frac{1}{2}$ since this will have no effect upon the expectation value:
@@
Violations of the Tsirelson inequality are manifestly associated with negativity in the Wigner function. A corollary is that a single coherent state, which has a completely positive Wigner function, will not violate the Tsirelson inequality; neither will a single energy eigenstate, for it is trivially of strictly odd or even parity. A neat proof, in terms of probability currents, that definite-parity states do not violate the Tsirelson inequality is provided in Section IV.1. Perhaps more surprising is that a superposition of at least three energy eigenstates is required to produce a violation. We provide a full proof of this fact in Appendix E.
@@ II.2 Maximally violating eigenstates
with an upper-bound Tsirelson violation $\langle A\rangle=1.1195$; or equivalently, by swapping the sign of the odd-parity eigenstate,
@@ II.3 Structure of the Tsirelson operator
Because the two normal operators commute, they cannot mix each other’s eigenspaces. The action of $U$ on the basis states $|n\rangle$ clearly partitions the Hilbert Space, $\mathcal{H}$, into 3 degenerate eigenspaces corresponding to its three eigenvalues $\lambda_{j}$:
@@
Figure 5: The Tsirelson operator in the energy eigenbasis with any finite energy cut-off can be interpreted as the adjacency matrix of a simple weighted graph. This is illustrated with a cut-off at $n=29$, where the nodes are energy eigenstates and the edges represent non-zero matrix elements of the Tsirelson operator. The colour represents the value of the matrix element.
@@ III.1 General Framework
We now come to the main issue, which is assessing the degree to which the dynamical assumptions behind the Tsirelson inequality hold. As discussed, like many other classical phenomena, uniform precession and its closely-related assumptions have no unique definition in quantum theory. We will therefore explore a number of inequivalent implementations which entail different measurement techniques to check it. Specifically, we will implement the following procedure in a number of different ways: